# COLT: Enhancing Video Large Language Models with Continual Tool Usage

### v2

@@ Introduction
Visual & Textual Embedding. Following (Lin et al. 2023a), we adopt LanguageBind (Zhu et al. 2023a) vision encoder ${g}(\cdot)$ to extract visual features. Then a linear projection layer ${f}_{\boldsymbol{\delta}}(\cdot)$ parameterized by $\boldsymbol{\delta}$ is applied to compress the visual information into LLM understandable feature space.
$\mathbf{H}_{\mathbf{v}}={f}_{\boldsymbol{\delta}}({g}(\mathbf{X}_{\mathbf{v}})),$ (1)
@@
where $\mathbf{H}_{\mathbf{v}}\in\mathbb{R}^{P\times C}$ denotes the visual embeddings, with $P$ and $C$ respectively representing the patch number and the feature dimension. For input user instructions $\mathbf{X}_{\mathbf{w}}$, we adopt the widely used BPE tokenizer (Sennrich, Haddow, and Birch 2016) to obtain the textual embeddings $\mathbf{H}_{\mathbf{w}}\in\mathbb{R}^{S\times C}$, where $S$ is the textual token number.
@@
Tool Codebook & Query Encoder. As shown in Figure 3, we set up a tool codebook $\mathbf{P}$ as the tool-specific memory for continual tool usage. The codebook consists of NN learnable tool prompts, i.e., $\mathbf{P}=\{\mathbf{P}_{n}\}_{n=1}^{N}$ with $\mathbf{P}_{n}\in\mathbb{R}^{C\times 1}$. For each input user instruction $\mathbf{X}_{\mathbf{q}}$, we aim to retrieve relevant tool prompts from the codebook. To achieve this, a query encoder ${f}_{\boldsymbol{\beta}}$ with parameters $\boldsymbol{\beta}$ is firstly employed to encode user instructions $\mathbf{X}_{\mathbf{q}}$ into $\mathbf{H}_{\mathbf{q}}\in\mathbb{R}^{C\times 1}$. Then, we compute the cosine similarities between the query embedding $\mathbf{H}_{\mathbf{q}}$ and each tool prompt $\mathbf{P}_{n}$, $n\in\{1,2,\cdots,N\}$. The tool prompts with top-$K$ highest similarity scores are selected.
@@
Training
The proposed COLT adopts a three-stage training methodology, with each stage featuring unique trainable parameters $\boldsymbol{\theta}$ and loss functions.
Stage 1: Video-to-text Alignment. This stage aims to train the projection layer ${f}_{\boldsymbol{\delta}}$ parameterized by $\boldsymbol{\delta}$, which acts as a visual tokenizer to align visual signals with pre-trained LLM word embedding. As shown in Figure 3, we freeze all the weights except the projection layer in this stage, i.e., trainable parameters $\boldsymbol{\theta}=\boldsymbol{\delta}$. Given the predicted response sequence $\mathbf{X}_{\mathbf{a}}=\{\mathbf{X}_{\mathbf{a}}^{1},\mathbf{X}_{\mathbf{a}}^{2},\cdots\mathbf{X}_{\mathbf{a}}^{L}\}$ of length $L$, we use the vanilla auto-regressive language modeling (LM) loss to supervise this stage of training:
$\mathcal{L}=\mathcal{L}_{\text{LM}}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\boldsymbol{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (3)
@@
Stage 2: Tool Codebook Pre-training. In this stage, we pre-train the tool codebook $\mathbf{P}$ and the query encoder ${f}_{\boldsymbol{\beta}}$ while keeping the other parts frozen, i.e., the trainable parameters are $\boldsymbol{\theta}=\{\mathbf{P},\boldsymbol{\beta}\}$. We firstly select the appropriate tool prompt $\mathbf{P}_{n^{*}}$ via Eq.(2). Then $\mathbf{P}_{n^{*}}$ serves as an additional condition for language modeling as follows:
$\mathcal{L}_{\text{LM}}^{\prime}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\boldsymbol{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},{\color[rgb]{0,0.6015625,0.3359375}\mathbf{P}_{n^{*}}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (4)
where $\mathbf{X}_{\mathbf{a}}^{i}$ and $\mathbf{X}_{\mathbf{a}}^{<i}$ is as defined in Eq. (3). However, directly optimizing $\mathcal{L}_{\text{LM}}^{\prime}$ will truncate the gradients w.r.t the query encoder ${f}_{\boldsymbol{\beta}}$ due to the non-derivable $\arg\operatorname{topk}$ operation in Eq. (2). Therefore, we resort to the straight-through estimator (Van Den Oord, Vinyals et al. 2017; Bengio, Léonard, and Courville 2013) to approximate the gradient computation and define the overall loss function as follows:
$\mathcal{L}=\mathcal{L}_{\text{LM}}^{\prime}+\lambda_{1}\underbrace{\left\|\operatorname{sg}\left[\mathbf{H}_{\mathbf{w}}\right]-\mathbf{P}_{n^{*}}\right\|_{2}^{2}}_{\mathcal{L}_{q}}+\lambda_{2}\underbrace{\left\|\mathbf{H}_{\mathbf{w}}-\operatorname{sg}[\mathbf{P}_{n^{*}}]\right\|_{2}^{2}}_{\mathcal{L}_{c}},$ (5)
@@
Stage 3: End-to-end Fine-tuning. We keep the vision encoder ${g}(\cdot)$ frozen and finetune remaining parts including the projection layer ${f}_{\boldsymbol{\delta}}$, tool codebook $\mathbf{P}$, textual encoder ${f}_{\boldsymbol{\beta}}$ and LLM parameterized by $\boldsymbol{\phi}$, i.e., the trainable parameters are $\boldsymbol{\theta}=\{\boldsymbol{\delta},\mathbf{P},\boldsymbol{\beta},\boldsymbol{\phi}\}$. The training loss is the same as Eq. (5).
Experiments
Experimental Settings
Training Datasets. In the first stage, we follow Video LLaVA (Lin et al. 2023a) to use LAION-$C$-SBU image subset (Schuhmann et al. 2021) and the filtered CC3M video dataset (Changpinyo et al. 2021) for video-to-text alignment. For the second and third stage, we use the combined dataset of 665K image-text instruction data from LLaVA v1.5 (Liu et al. 2023a) and 100K video-text instruction data from Video-ChatGPT (Maaz et al. 2023). The tool-use instruction data of our collected VideoToolBench is introduced into continual training sequentially. To ensure conformity in data formatting, we reformat all the instruction tuning datasets to the thought-action-value pattern. Refer to supplementary materials for more details.
Table 2: Comparisons with state-of-the-art methods on zero-shot video-question answering. “Acc.” denotes accuracy (%) and “Score” denotes the relative score from 0 to 5 assigned by GPT (Brown et al. 2020). The best performance is in bold and the second best is underlined. The backend LLMs include Vicuna-7B (Chiang et al. 2023) and LLaMA-7B (Touvron et al. 2023a).
@@
Video-LLaMA (Zhang, Li, and Bing 2023) Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1
VideoChat (Li et al. 2023b) Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2
Video-ChatGPT (Maaz et al. 2023) Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter (Liu et al. 2023b) Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2
LLaMA-VID (Li, Wang, and Jia 2023) Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3
LLaMA-VID (Li, Wang, and Jia 2023) Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3
Video-LLaVA (Lin et al. 2023a) Vicuna-7B 70.7 3.9 59.2 3.5 45.3 3.3
Chat-UniVi (Jin et al. 2023) Vicuna-7B 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-Adapter (Zhang et al. 2023b) LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7
VideoChat2 (Li et al. 2023c) Vicuna-7B 70.0 3.9 54.1 3.3 49.1 3.3
ST-LLM (Liu et al. 2024b) Vicuna-7B 74.6 3.9 63.2 3.4 50.9 3.3
@@
Implementation Details. For vision encoder $\boldsymbol{g}(\cdot)$, we choose pre-trained LanguageBind (Zhu et al. 2023a) with ViT-L/14 (Dosovitskiy et al. 2020). The text tokenizer is derived from LLaMA (Touvron et al. 2023a) with a vocabulary size of 32,000 and Vicuna-7B v1.5 (Chiang et al. 2023) is employed as the large language model. We uniformly sample 8 frames from each video and each frame is resized to 224 $\times$ 224. We set the batch size to 256 for the first stage and 128 for the second and third stage. AdamW optimizer is used with a cosine decay schedule. We set learning rate to $\times 10^{-4}, 1×1$$\times 10^{-4}$, and $1\times 10^{-5}$ for three stage training, respectively. The balancing weight $\lambda_{1}$ and $\lambda_{2}$ in Eq. (5) are set to 1 and 0.25. The codebook size NN and selected prompt number $K$ is set to 50 and 3, respectively.
Comparisons with Video LLMs
@@
InstructBLIP (Dai et al. 2024) 32.5 20.0 16.5 46.0 24.5 46.0 51.0 26.0 37.5 22.0 23.0
LLaVA (Liu et al. 2024a) 36.0 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5
VideoChatGPT (Maaz et al. 2023) 32.7 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0
VideoLLaMA (Zhang, Li, and Bing 2023) 34.1 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5
VideoChat (Li et al. 2023b) 35.5 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0
VideoChat2text (Li et al. 2023c) 34.7 24.5 27.0 49.5 27.0 38.0 53.0 28.0 40.0 25.5 27.0
VideoChat2 (Li et al. 2023c) 51.1 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0
GPT-4V (OpenAI 2023) 43.5 55.5 63.5 72.0 46.5 73.5 18.5 59.0 29.5 12.0 40.5
@@
InstructBLIP (Dai et al. 2024) 32.5 46.5 42.5 26.5 40.5 32.0 25.5 30.0 25.5 30.5 38.0
LLaVA (Liu et al. 2024a) 36.0 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0
VideoChatGPT (Maaz et al. 2023) 32.7 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5
VideoLLaMA (Zhang, Li, and Bing 2023) 34.1 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0
VideoChat (Li et al. 2023b) 35.5 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0
VideoChat2text (Li et al. 2023c) 34.7 38.5 41.5 27.5 32.5 46.5 26.5 36.0 33.0 32.0 40.0
VideoChat2 (Li et al. 2023c) 51.1 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5
GPT-4V (OpenAI 2023) 43.5 83.5 39.0 12.0 22.5 45.0 47.5 52.0 31.0 59.0 11.0
@@
Table 5: Comparisons with SOTA methods on test split of VideoToolBench. AR, VRD, and ASR denote action recognition, video relation detection and automatic speech recognition, respectively. ”Acc.” denotes accuracy (%) and ”Score” denotes the GPT-evaluated score.
@@
[…]
Evaluation Benchmarks. Our experiments are carried out on both established video LLM benchmarks and self-built tool-using datasets: 1) zero-shot video-question answering. We experiment on commonly used open-ended question-answer datasets: MSVD-QA (Chen and Dolan 2011), MSRVTT-QA (Xu et al. 2016), and ActivityNet-QA (Yu et al. 2019). Following (Li et al. 2023b; Lin et al. 2023a), we use GPT-assisted evaluation to assess the model’s capabilities by reporting the accuracy and relative score; 2) MVBench (Li et al. 2023c). This benchmark consists of 20 demanding video tasks, each comprising 200 samples presented as multiple-choice questions. These tasks offer a thorough and unbiased evaluation of a model’s capacity to comprehend videos. We report the choice accuracy as the metric; 3) VideoToolBench. We build this dataset to probe the abilities enabled by tool proficiency. Since most existing video LLMs only support textual outputs, we select three tools (i.e., action recognition, video relation detection, and automatic speech recognition) and compare our COLT to state-of-the-art video LLMs (Lin et al. 2023a; Li et al. 2023c, b). The GPT-evaluated accuracy and scores are reported.
Performance Analysis. The comparison results on zero-shot video-question answering, MVBench, and VideoToolBench are summarized in Table 2, Table 3, and Table 5, respectively. We can conclude with the following findings. 1) Both COLT $_{\text{joint}}$ and COLT $_{\text{5$\times$2}}$ demonstrate superior performance compared to preceding state-of-the-art methodologies. For example on MSRVTT-QA (c.f. Table 2), COLT $_{\text{joint}}$ remarkably surpasses the previous best performing method VideoChat2 (Li et al. 2023c) by 8.2% on the metric of accuracy; 2) On the test set of VideoToolBench (c.f. Table 5), our COLT significantly outperforms previous methods by a remarkable margin, which demonstrates that our method is more effective in scenarios that require invoking tools to complete; 3) Even with a rather lengthy learning curve, COLT $_{\text{10$\times$1}}$ still achieves comparable performance with previous SOTA methods; 4) The performance of COLT $_{\text{10$\times$1}}$ is slightly worse than COLT $_{\text{5$\times$2}}$, which is consistent with the intuition that tool learning becomes increasingly challenging when facing a lengthening tool-chain.
Comparisons with Continual Learning Methods
Table 6: Comparisons (%) with continual learning methods VideoToolBench. AA and AF denote average accuracy and average forgetting, respectively.
@@
L2P (Wang et al. 2022c) 72.4 8.3 65.7 12.4
Dual (Wang et al. 2022b) 75.1 6.9 68.6 7.5
CODA (Smith et al. 2023) 76.3 7.0 71.0 7.2

### v3

@@ 4.1 Architecture
Visual & Textual Embedding. Following (lin2023video), we adopt LanguageBind vision encoder (zhu2023languagebind) ${g}(\cdot)$ to extract visual features. Then, a linear projection layer ${f}_{\bm{\delta}}(\cdot)$ parameterized by $\bm{\delta}$ is applied to compress the visual information into an LLM understandable feature space.
$\mathbf{H}_{\mathbf{v}}={f}_{\bm{\delta}}({g}(\mathbf{X}_{\mathbf{v}})),$ (1)
@@
where $\mathbf{H}_{\mathbf{v}}\in\mathbb{R}^{P\times C}$ denotes the visual embeddings, with $P$ and $C$ respectively representing the patch number and the feature dimension. For input user instructions $\mathbf{X}_{\mathbf{w}}$, we adopt the widely used BPE tokenizer (sennrich2016neural) to obtain the textual embeddings $\mathbf{H}_{\mathbf{w}}\in\mathbb{R}^{S\times C}$, where $S$ is the textual token number.
@@
Tool Codebook & Query Encoder. As shown in Figure 3, we set up a tool codebook $\mathbf{P}$ as the tool-specific memory for continual tool usage. The codebook consists of $N$ learnable tool prompts, i.e., $\mathbf{P}=\{\mathbf{P}_{n}\}_{n=1}^{N}$ with $\mathbf{P}_{n}\in\mathbb{R}^{C\times 1}$. For each input user instruction $\mathbf{X}_{\mathbf{q}}$, we aim to retrieve relevant tool prompts from the codebook. To achieve this, a query encoder ${f}_{\bm{\beta}}$ with parameters $\bm{\beta}$ is firstly employed to encode user instructions $\mathbf{X}_{\mathbf{q}}$ into $\mathbf{H}_{\mathbf{q}}\in\mathbb{R}^{C\times 1}$. Then, we compute the cosine similarities between the query embedding $\mathbf{H}_{\mathbf{q}}$ and each tool prompt $\mathbf{P}_{n}$, $n\in\{1,2,\cdots,N\}$. The tool prompts with the top-$K$ highest similarity scores are selected.
@@
4.2 Training
The proposed COLT adopts a three-stage training methodology, with each stage featuring unique trainable parameters $\bm{\theta}$ and loss functions.
Stage 1: Video-to-text Alignment. This stage aims to train the projection layer ${f}_{\bm{\delta}}$ parameterized by $\bm{\delta}$, which acts as a visual tokenizer to align visual signals with pre-trained LLM word embedding. As shown in Figure 3, we freeze all the weights except the projection layer in this stage, i.e., trainable parameters $\bm{\theta}=\bm{\delta}$. Given the predicted response sequence $\mathbf{X}_{\mathbf{a}}=\{\mathbf{X}_{\mathbf{a}}^{1},\mathbf{X}_{\mathbf{a}}^{2},\cdots\mathbf{X}_{\mathbf{a}}^{L}\}$ of length $L$, we use the vanilla auto-regressive language modeling (LM) loss to supervise this stage of training:
$\mathcal{L}=\mathcal{L}_{\text{LM}}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\bm{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (3)
@@ 4.2 Training
Stage 2: Tool Codebook Pre-training. In this stage, we pre-train the tool codebook $\mathbf{P}$ and the query encoder ${f}_{\bm{\beta}}$ while keeping the other parts frozen, i.e., the trainable parameters are $\bm{\theta}=\{\mathbf{P},\bm{\beta}\}$. We firstly select the appropriate tool prompt $\mathbf{P}_{n^{*}}$ via Eq.equation 2. Then $\mathbf{P}_{n^{*}}$ serves as an additional condition for language modeling as follows:
$\small\mathcal{L}_{\text{LM}}^{\prime}=-\frac{1}{L}\sum_{i=1}^{L}\;\log p_{\bm{\theta}}(\mathbf{X}_{\mathbf{a}}^{i}\mid\mathbf{H}_{\mathbf{v}},\mathbf{H}_{\mathbf{w}},{\color[rgb]{0,0.6015625,0.3359375}\mathbf{P}_{n^{*}}},\mathbf{X}_{\mathbf{a}}^{<i}),$ (4)
where $\mathbf{X}_{\mathbf{a}}^{i}$ and $\mathbf{X}_{\mathbf{a}}^{<i}$ is as defined in Eq. equation 3. However, directly optimizing $\mathcal{L}_{\text{LM}}^{\prime}$ will truncate the gradients w.r.t the query encoder ${f}_{\bm{\beta}}$ due to the non-derivable $\arg\operatorname{topk}$ operation in Eq. equation 2. Therefore, we resort to the straight-through estimator (van2017neural; bengio2013estimating) to approximate the gradient computation and define the overall loss function as follows:
$\small\mathcal{L}=\mathcal{L}_{\text{LM}}^{\prime}+\lambda_{1}\underbrace{\left\|\operatorname{sg}\left[\mathbf{H}_{\mathbf{w}}\right]-\mathbf{P}_{n^{*}}\right\|_{2}^{2}}_{\mathcal{L}_{q}}+\lambda_{2}\underbrace{\left\|\mathbf{H}_{\mathbf{w}}-\operatorname{sg}[\mathbf{P}_{n^{*}}]\right\|_{2}^{2}}_{\mathcal{L}_{c}},$ (5)
@@
Stage 3: End-to-end Fine-tuning. We keep the vision encoder ${g}(\cdot)$ frozen and finetune remaining parts including the projection layer ${f}_{\bm{\delta}}$, tool codebook $\mathbf{P}$, textual encoder ${f}_{\bm{\beta}}$ and LLM parameterized by $\bm{\phi}$, i.e., the trainable parameters are $\bm{\theta}=\{\bm{\delta},\mathbf{P},\bm{\beta},\bm{\phi}\}$. The training loss is the same as Eq. equation 5.
5 Experiments
5.1 Experimental Settings
Training Datasets. In the first stage, we follow Video LLaVA lin2023video to use LAION-CC-SBU image subset schuhmann2021laion and the filtered CC3M video dataset changpinyo2021conceptual for video-to-text alignment. For the second and third stage, we use the combined dataset of 665K image-text instruction data from LLaVA v1.5 liu2023improved and 100K video-text instruction data from Video-ChatGPT maaz2023video. The tool-use instruction data of our collected VideoTool is introduced into continual training sequentially. To ensure conformity in data formatting, we reformat all the instruction tuning datasets to the thought-action-value pattern. Refer to supplementary materials for more details.
Table 2: Comparisons with state-of-the-art methods on zero-shot video-question answering. “Acc." denotes accuracy (%) and “Score" denotes the relative score from 0 to 5 assigned by GPT brown2020language. The best performance is in bold and the second best is underlined. The backend LLMs include Vicuna-7B vicuna2023 and LLaMA-7B touvron2023llama.
@@ 5.1 Experimental Settings
Video-LLaMA zhang2023video Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1
VideoChat li2023videochat Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2
Video-ChatGPT maaz2023video Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter liu2023one Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2
LLaMA-VID li2023llama Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3
LLaMA-VID li2023llama Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3
Video-LLaVA lin2023video Vicuna-7B 70.7 3.9 59.2 3.5 45.3 3.3
Chat-UniVi jin2023chat Vicuna-7B 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-Adapter zhang2023llama LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7
VideoChat2 li2023mvbench Vicuna-7B 70.0 3.9 54.1 3.3 49.1 3.3
ST-LLM liu2024st Vicuna-7B 74.6 3.9 63.2 3.4 50.9 3.3
@@
Implementation Details. For vision encoder $\bm{g}(\cdot)$, we choose pre-trained LanguageBind zhu2023languagebind with ViT-L/14 dosovitskiy2020image. The text tokenizer is derived from LLaMA touvron2023llama with a vocabulary size of 32,000, and Vicuna-7B v1.5 vicuna2023 is employed as the large language model. We uniformly sample 8 frames from each video, and each frame is resized to 224 $\times$ 224. We set the batch size to 256 for the first stage and 128 for the second and third stages. AdamW optimizer is used with a cosine decay schedule. We set learning rate to $\times 10^{-4}, 1×1$$\times 10^{-4}$, and $1\times 10^{-5}$ for three stage training, respectively. The balancing weight $\lambda_{1}$ and $\lambda_{2}$ in Eq. equation 5 are set to 1 and 0.25. The codebook size $N$ and selected prompt number $K$ are set to 50 and 3, respectively. COLT is trained on 8 NVIDIA A100 GPUs (80 GB memory each), and the full training process takes approximately 10 hours.
Evaluation Metrics of Tool Continual Learning. We detail the metrics of average accuracy and average forgetting used in Sec. 5.3.
$\text{AA}_{k}=\frac{1}{k}\sum_{j=1}^{k}a_{k,j}.$ (6)
$f_{j}^{k}=\max_{l\in\{1,\cdots,k-1\}}a_{l,j}-a_{k,j},\quad\forall j<k.$ (7)
$\text{AF}_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}f_{j}^{k}.$ (8)
5.2 Comparisons with Video LLMs
@@
InstructBLIP dai2024instructblip 32.5 20.0 16.5 46.0 24.5 46.0 51.0 26.0 37.5 22.0 23.0
LLaVA liu2024visual 36.0 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5
VideoChatGPT maaz2023video 32.7 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0
VideoLLaMA zhang2023video 34.1 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5
VideoChat li2023videochat 35.5 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0
VideoChat2text li2023mvbench 34.7 24.5 27.0 49.5 27.0 38.0 53.0 28.0 40.0 25.5 27.0
VideoChat2 li2023mvbench 51.1 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0
GPT-4V gpt4v 43.5 55.5 63.5 72.0 46.5 73.5 18.5 59.0 29.5 12.0 40.5
@@
InstructBLIP dai2024instructblip 32.5 46.5 42.5 26.5 40.5 32.0 25.5 30.0 25.5 30.5 38.0
LLaVA liu2024visual 36.0 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0
VideoChatGPT maaz2023video 32.7 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5
VideoLLaMA zhang2023video 34.1 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0
VideoChat li2023videochat 35.5 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0
VideoChat2text li2023mvbench 34.7 38.5 41.5 27.5 32.5 46.5 26.5 36.0 33.0 32.0 40.0
VideoChat2 li2023mvbench 51.1 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5
GPT-4V gpt4v 43.5 83.5 39.0 12.0 22.5 45.0 47.5 52.0 31.0 59.0 11.0
@@
(a) Comparisons with SOTA methods on test split of VideoTool. AR, VRD, and ASR denote action recognition, video relation detection and automatic speech recognition, respectively. "Acc." denotes accuracy (%) and "Score" denotes the GPT-evaluated score.
@@
(b) Comparisons (%) with continual learning methods on VideoTool. AA and AF denote average accuracy and average forgetting, respectively.
@@
L2P wang2022learning 72.4 8.3 65.7 12.4
Dual wang2022dualprompt 75.1 6.9 68.6 7.5
CODA smith2023coda 76.3 7.0 71.0 7.2
@@
[…]
Evaluation Benchmarks. Our experiments are carried out on both established video LLM benchmarks and self-built tool-using datasets: 1) zero-shot video-question answering. We experiment on commonly used open-ended question-answer datasets: MSVD-QA chen2011collecting, MSRVTT-QA xu2016msr, and ActivityNet-QA yu2019activitynet. Following li2023videochat; lin2023video, we use GPT-assisted evaluation to assess the model’s capabilities by reporting the accuracy and relative score; 2) MVBench li2023mvbench. This benchmark consists of 20 demanding video tasks, each comprising 200 samples presented as multiple-choice questions. These tasks offer a thorough and unbiased evaluation of a model’s capacity to comprehend videos. We report the choice accuracy as the metric; 3) VideoTool. We built this dataset to probe the abilities enabled by tool proficiency. Since most existing video LLMs only support textual outputs, we select three tools (i.e., action recognition, video relation detection, and automatic speech recognition) and compare our COLT to state-of-the-art video LLMs lin2023video; li2023mvbench; li2023videochat. The GPT-evaluated accuracy and scores are reported.
Performance Analysis. The comparison results on zero-shot video-question answering, MVBench, and VideoTool test split are summarized in Table 2, Table 3, and Table 5(a), respectively. We can conclude with the following findings. 1) Both COLT $_{\text{joint}}$ and COLT $_{\text{5$\times$2}}$ demonstrate superior performance compared to preceding state-of-the-art methodologies. For example on MSRVTT-QA (c.f. Table 2), COLT $_{\text{joint}}$ remarkably surpasses the previous best performing method VideoChat2 li2023mvbench by 8.2% on the metric of accuracy; 2) On the test set of VideoToolBench (c.f. Table 5(a)), our COLT consistently outperforms prior approaches by a clear margin, including both existing video-VQA models and representative tool-based agents, demonstrating its effectiveness in tool-intensive scenarios; 3) Even with a rather lengthy learning curve, COLT $_{\text{10$\times$1}}$ still achieves comparable performance with previous SOTA methods; 4) The performance of COLT $_{\text{10$\times$1}}$ is slightly worse than COLT $_{\text{5$\times$2}}$, which is consistent with the intuition that tool learning becomes increasingly challenging when facing a lengthening tool-chain.
5.3 Comparisons with Continual Learning Methods